1. A) How many grams of Mg(OH)2 will be needed to neutralize 25mL of stomach acid...

1.

A) How many grams of Mg(OH)2 will be needed to neutralize 25mL of stomach acid if stomach acid is 0.10 M HCl?

B) How many mL of a 0.10 M NaOH solution are needed to neutralize 15 mL of 0.20 M H3PO4 solution?
Answer #1

A)

Balanced chemical equation is:

2 HCl + Mg(OH)2 ---> MgCl2 + 2 H2O

lets calculate the mol of HCl

volume , V = 25 mL

= 2.5*10^-2 L

use:

number of mol,

n = Molarity * Volume

= 0.1*2.5*10^-2

= 2.5*10^-3 mol

According to balanced equation

mol of Mg(OH)2 reacted = (1/2)* moles of HCl

= (1/2)*2.5*10^-3

= 1.25*10^-3 mol

This is number of moles of Mg(OH)2

Molar mass of Mg(OH)2,

MM = 1*MM(Mg) + 2*MM(O) + 2*MM(H)

= 1*24.31 + 2*16.0 + 2*1.008

= 58.326 g/mol

use:

mass of Mg(OH)2,

m = number of mol * molar mass

= 1.25*10^-3 mol * 58.33 g/mol

= 7.291*10^-2 g

Answer: 7.29*10^-2 g

B)

Balanced chemical equation is:

H3PO4 + 3 NaOH ---> Na3PO4 + 3 H2O

Here:

M(H3PO4)=0.2 M

M(NaOH)=0.1 M

V(H3PO4)=15.0 mL

According to balanced reaction:

3*number of mol of H3PO4 =1*number of mol of NaOH

3*M(H3PO4)*V(H3PO4) =1*M(NaOH)*V(NaOH)

3*0.2 M *15.0 mL = 1*0.1M *V(NaOH)

V(NaOH) = 90 mL

Answer: 90 mL